1. Calculate the equilibrium level of income if C = R100 million + 0,8Y and Ī = R125 million.[5
1
Expert's answer
2019-08-05T11:46:19-0400

Equilibrium condition is

"S=I"

"s=Y-C=Y-a-b\times{Y}=-a+(1-b)\times{Y}"

for

"I=I"

The condition becomes

"-a+(1-b)\times{Y}=I"

Therefore the equilibrium level of income is

"Y=\frac{1}{1-b\times(a+I)}"

From the consumption equation

"C=100+0.8\times{Y}"

We find a and 1-b


"-a=-100""b=1-0.8=0.2"


Substituting into saving and investment equation gives

"-100+0.2\times{Y}=125"

"0.2\times{Y}=225""Y=Rs 1125m"